OpenStudy (anonymous):

Write the standard form of the equation of each line passing through P with slope m. P=(10,2); m=1/1

OpenStudy (anonymous):

y=m*x+c where m= slope, c is a constant called the y intercept Since, 10,2 lies on the line, so it must satisfy the above equation. Or, 2=1/1*10+c or c=1/5 So the equation is y=m*x+1/5 IS the slope 1/1?? I mean, it's kind of awkward 'cause it means just 1.

OpenStudy (anonymous):

yes it is 1/1 and 1/1 =1

OpenStudy (anonymous):

so what do we do next?

OpenStudy (anonymous):

Wait. I made a mistake. c= 2-10= -8 So the equation is y=x-8 Or, you could write the above as x/8-y/8=1

OpenStudy (anonymous):

ok thanks